Category Archive: News

General announcements

SDSC Planned Maintenance – East Data Center emergency power – 28 Jan 2017, 0630

This maintenance has been completed, and power has been restored.  Please let us know if there are any further issues via normal support channels. Best, –Jeff         Good morning, We are currently within this maintenance window.  All emergency power PDUs are offline in the East data center, and all HVAC systems are …

Continue reading »

SDSC Outage Notification – Oracle services – 19 Jan 2017, 19:50

[Update : 20:56] The cause was found. During routine maintenance an assumed-unused package was removed. Once the package was re-added, Oracle began working again. Services are up — please contact support for any continuing issues. — At 19:50 on 19 Jan 2017 the primary Oracle nodes serving the SDSC Footprints system and SAM/ART accounting started …

Continue reading »

SDSC Planned Maintenance – linux patch/reboot – 31 January 2017, 17:00-23:00

On Tuesday, 31 January 2016 between 17:00 and 23:00, general system maintenance, including patching and reboots, will occur for the following systems. Web sites and services hosted on the below systems will be affected. Systems will be unavailable during reboots. Please save all work prior to the 5pm maintenance period. Affected Linux systems: bastiat batman …

Continue reading »

SDSC Planned Maintenance – network time protocol (NTP) – 9 January 2017

SDSC has made changes to prevent access to incorrect NTP servers. Please remember that local NTP servers to use within the SDSC datacenter are: ns0.sdsc.edu (132.249.20.25) ns1.sdsc.edu (198.202.75.26) moe.sdsc.edu (132.249.20.63) Please contact service providers with questions or problems following this modification.

SDSC Outage Notification – limited Linux guest outage – 14 Dec 2016, 19:45

[Update, 00:05, 15 Dec 2016] guest ‘rupee’ is now online. [Update, 23:37] All guests except ‘rupee’ are online. Please contact support with any issues or concerns [Update, 22:05] Engineers replaced the faulty disk controller and verified the guests function. Guests were then shut down again so that engineers can perform a reboot and ensure the …

Continue reading »

SDSC Planned Maintenance – Linux patch/reboot – 20 Dec 2016, 17:00-23:00

[Update, 20 Dec 2016, 23:08] All maintenance is complete. Please contact the support team with any questions or concerns —– Colleagues: On Tuesday, 20 December 2016 between 17:00 and 23:00, general system maintenance, including patching and reboots, will occur for the following systems. Systems will be unavailable during reboots. Please save all work prior to …

Continue reading »

SDSC Outage Notification – CDS2 / West UPS Power – 3 Nov 2016, 10:40

[Update – 20:14] The UPS has been reactivated and is protecting the systems. —– At approximately 10:40 the CDS2 UPS system lost power due to a short circuit condition. The power was immediately restored with the UPS system being temporarily bypassed. Updates will be forthcoming. The scope of the outage would be computer which were …

Continue reading »

SDSC Emergency Maintenance – Linux patch/reboot – 1 Nov 2016, 20:00-23:00

[Update – 20:41] All patches and reboots have been applied. —- SDSC will be applying critical patches to the Linux environment tonight, 1 Nov 2016 starting at 8pm. This maintenance will require a reboot of all systems listed below. Please contact support@sdsc.edu with any questions or concerns. Updates will be posted to http://status.sdsc.edu as the …

Continue reading »

SDSC Emergency Maintenance – Linux patch/reboot – 27 Oct 2016, 20:00-23:00

[Update – 22:04] Updates and reboots have been completed. Please contact support with any questions or problems —- SDSC will be applying critical patches to the Linux environment tomorrow night, 27 Oct 2016 starting at 8pm. This maintenance will require a reboot of all systems listed below. Please contact support@sdsc.edu with any questions or concerns. …

Continue reading »

SDSC Planned Maintenance – core NFS server replacement – 5 Nov 2016, 08:00-13:00

[Update, 15:10] All systems are up and services restored. Please contact support with any quesitons or problems. [Update, 14:02] Permissions problems were solved. Due to the migration of fileserver from Solaris to Linux, some default ownership changed needed to new values. Also, some networking changes needed to be changed. [Update, 13:10] Many systems are back …

Continue reading »

Older posts «

» Newer posts